OpenStudy (anonymous):

solve the system by substitution -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (nincompoop):

what does substitution mean?

OpenStudy (anonymous):

.....

OpenStudy (nincompoop):

do you know what it means?

OpenStudy (anonymous):

Please don't bother helping me if you don't know that..

OpenStudy (nincompoop):

I don't know it? I am asking you if you knew it I wouldn't bother helping if I didn't know anything

OpenStudy (nincompoop):

but I guess you need help beyond math, because your attitude seems to be a problem too

OpenStudy (anonymous):

Jesus, go away.

OpenStudy (nincompoop):

Jesus is not my name :P

OpenStudy (anonymous):

How old are you, 12?

OpenStudy (anonymous):

girls, stop fighting, you're both pretty. now can I go home?

OpenStudy (anonymous):

emily - solve the first equation for y and the third equation for z. use the expressions you find for y and z to solve for x in the second equation. have fun

OpenStudy (anonymous):

Can you show me step by step?

OpenStudy (anonymous):

I just told you the steps. showing would be doing the work for you.

OpenStudy (anonymous):

I can show you a simpler case. x-y=0 x+y=2 a 2x2 system of equations solve the first equation for x x=y use the expression for x in the second equation to make the second equation have only one variable y+y=2 solve for y 2y=2, y=1 then solve for x x=y=1

OpenStudy (anonymous):

have fun

OpenStudy (anonymous):

so, it's 2y=2? Then x=y=1? I dunno what to do next.

OpenStudy (anonymous):

I was showing the process in a simpler case, nothing to do with your problem. follow the steps I outlined earlier to solve your problem.

OpenStudy (anonymous):

please tell me what to do. I'm desperate. :(

OpenStudy (nincompoop):

so do you know what it means to substitute?

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

solve the first equation for y and the third equation for z. use the expressions you find for y and z to solve for x in the second equation.

OpenStudy (anonymous):

You aren't very helpful @Peter14

OpenStudy (anonymous):

solve the first equation for y y=(something) solve the third equation for z z=(someotherthing) use the expressions for y and z in the second equation to solve for x -4x+4y+5z=7 -4x+4(something)+5(someotherthing)=7 (something) will have a z-term in it too, so put (someotherthing) in its place, then solve for x.

OpenStudy (anonymous):

...you could have just told me you wanted me to do it for you at the beginning.

OpenStudy (usukidoll):

oh no they were helpful...it's you who is being rude. ;) And by the way, Matrix style for the win!

OpenStudy (usukidoll):

2 x 3 augmented matrix but heyyyy if it's solve by substitution, it means algebra style and that's tedious :D

OpenStudy (anonymous):

How on earth am I being rude?

OpenStudy (anonymous):

Asking for help? Stop picking on me for wanting fluttering help. Jesus Christ.

OpenStudy (usukidoll):

calling nin Jesus is one which is using the lord's name in vain. Two we help users, not spoon feed answers.

OpenStudy (usukidoll):

whatever I spent 7 years on Mathematics...so >:P

OpenStudy (anonymous):

How old are you?

OpenStudy (usukidoll):

none of your business

OpenStudy (anonymous):

UsukiDoll, this post is no longer worth our time.

OpenStudy (anonymous):

Hm. Must be old.

OpenStudy (usukidoll):

duh! And yeah let's go @Peter14

OpenStudy (anonymous):

So, you're old, with a my Little Pony picture set as your profile picture, on this site helping kids at midnight? 0_ o

OpenStudy (usukidoll):

And you're being disrespectful beyond recognition.

OpenStudy (usukidoll):

so what there are bronies out there with the my little pony craze

OpenStudy (anonymous):

No, asking for help. Just getting pissed cause of people like you

OpenStudy (usukidoll):

aren't you in connections academy?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

you aren't asking for help anymore, I gave you all the help you need. I stopped one step short of doing the problem for you.

OpenStudy (usukidoll):

what Peter said

OpenStudy (anonymous):

stop. get up and get a drink of water. look at my explanation. then do the math.

OpenStudy (anonymous):

you can do it.

OpenStudy (anonymous):

No. You didn't do pellet, peter pan

OpenStudy (anonymous):

peter rabbit, get it right!

OpenStudy (usukidoll):

oh yes he did. It's you who is not doing the problem

OpenStudy (anonymous):

Oh, yes. That's it xD

OpenStudy (anonymous):

solve the first equation for y y=(something) solve the third equation for z z=(someotherthing) use the expressions for y and z in the second equation to solve for x -4x+4y+5z=7 -4x+4(something)+5(someotherthing)=7 (something) will have a z-term in it too, so put (someotherthing) in its place, then solve for x. good luck :)

OpenStudy (nincompoop):

even if she learned that portion of math I doubt people will have patience and even consideration to deal with a person who is both ingrate and bad-mannered